Given that f is the real valued function defined by f(x) = 1 / (1 + e^x). And I = (1/4, 1/2). We define the sequence (Un) by Uo = 1/4 and for all n E N where N = set of natural numbers, and E represents "an element of", U(n +1) = f(Un).

(i)Show that Un E I.
(ii) Show that for all n E N, we have ㅣU(n+1) - sㅣ <= 1/4ㅣUn - sㅣ Where s Eb(1/4, 1/2)
1
Expert's answer
2021-02-02T05:46:20-0500

"\displaystyle
(i)\\ f(x) = \frac{1}{1 + e^x}\\

f(U_n) = \frac{1}{1 + e^{U_n}}\\


\frac{2}{e^x + e^{-x}} = \frac{1}{\cosh{x}}\\

\frac{1}{1 + e^{2x}} = \frac{e^{-x}}{2\cosh{x}}\\

f(x) = \frac{1}{1 + e^{x}} = \frac{e^{-\frac{x}{2}}}{2\cosh\left(\frac{x}{2}\right)}\\

f(0) = \frac{1}{1 + e^{0}} = \frac{1}{2}\\

\textsf{If}\,\, x \in I = \left(\frac{1}{4}, \frac{1}{2}\right) \,\, \textsf{then} \\ 

f(x) < f(0) = \frac{1}{2}, \\

f(x) > f(1) = \frac{1}{1 + e} > \frac{1}{1 + 3} = \frac{1}{4},\\

\textsf{and hence,}\,\, f(x) \in I. \\

u_0 \in I \,\, \textsf{implies}\,\, u_1 = f(u_0) \in I. \\


u_n \in I\,\, \textsf{for} \,\, n \in \mathbb{N} \textsf{implies}\\

u_{n + 1} = f(u_n) \in I\,\, \textsf{and, by induction,} \\

u_n \in I \,\, \forall \,\, n \in \mathbb{N} \\
\textbf{\textsf{Proved}}\\
(ii) \\

h(x) = f(x) - x = \frac{1}{1 + e^x} - x\,\,\\
 \textsf{is a continuous function.}\\

h\left(\frac{1}{4}\right) = f\left(\frac{1}{4}\right) - \frac{1}{4} > f(1) - \frac{1}{4} \\=
\frac{1}{1 + e} - \frac{1}{4} > \frac{1}{1 + 3} - \frac{1}{4} = 0 \\

h\left(\frac{1}{2}\right) = f\left(\frac{1}{2}\right) - \frac{1}{2} < f(0) - \frac{1}{2} = 0\\

\textsf{Therefore, by the intermediate value}\\
\textsf{theorem we have that there exists a root}\,\, s \in I \\
\textsf{of this function, that is, a number}\,\, s \,\, \textsf{such that}\\


f(s) = s\\

\textsf{By MVT}\\

\frac{u_{n + 1} - s}{u_n - s} = \frac{f(u_n) - f(s)}{u_n - s} = f'(\epsilon_n) \\

\textsf{for} \,\,\epsilon_n \in (u_n, s) \\


1 + e^x \geq 2e^{\frac{x}{2}} \,\,\,\,\, (AP \geq GP) \\

|f'(\epsilon_n)| = \frac{e^x}{(1 + e^x)^2} \leq \frac{e^x}{(2e^{\frac{x}{2}})^2} \leq \frac{e^x}{4e^x} = \frac{1}{4} \\


\textsf{Comparing the last two formulae, we have}\\
|u_{n + 1} - s| \leq \frac{|u_n - s|}{4}, \\
\textbf{\textsf{Proved}}"